OpenStudy (anonymous):

please help me The measure of the largest angle of a triangle is 90◦ more than the measure of the smallest angle, and the measure of the remaining angle is 30◦ more than the measure of the smallest angle. Find the measure of each angle. Show your work in obtaining your answer.

OpenStudy (anonymous):

can you write equation to model this?

OpenStudy (anonymous):

i dont know how

OpenStudy (anonymous):

smallest = s biggest = b medium = m b = s + 90 m = s + 30 s + m + b = 180 s + s + 30 + s + 90 = 180 3 s = 60 etc.....

OpenStudy (anonymous):

what am i supposed to solve what equation

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Do you want the equation or the answer?

OpenStudy (anonymous):

i want the equation that im supposed to be solving in order to get the answer

OpenStudy (anonymous):

is it system of equations?

OpenStudy (anonymous):

I don't remember the exact equation, just what I wrote which is what my teacher had written down for me. The answers are S=60 M=30 B=90 that's the best I can do. Sorry

OpenStudy (anonymous):

oh okay, thanks anyway

OpenStudy (anonymous):

i still need help if you all know

OpenStudy (anonymous):

what am i supposed to e solving

OpenStudy (anonymous):

x+(x+30)+(x+90)=180 3x+120=180 3x=180-120=60 x=?

OpenStudy (anonymous):

20

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

now find the other angles.

OpenStudy (anonymous):

x+30=? x+90=?

OpenStudy (anonymous):

50 and 110?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

can you help me with another problem please?

OpenStudy (anonymous):

sure,i will try.

OpenStudy (anonymous):

do i post it here or new post?

OpenStudy (anonymous):

anywhere

OpenStudy (anonymous):

In each of the following systems of equations, look at the system and determine whether it has infinitely many solutions, no solution, or one solution. Explain in each case how you came to your conclusion. You should not attempt to solve any of the systems. a) Let k represent a Real Number. y = 2x + k y = 2x − k b) Let a represent a Real Number. y = −4x + a y= 1/4x-a c) Let c represent a Real Number. 2x + 6y = 4c 3x + 9y = 6c

OpenStudy (anonymous):

let us solve one by one 1 y=2x+k y=2x-k 2x+k=2x-k k=-k k+k=0 2k=0 k=0 so if k=0 y=2x has infinite solutions case 2.if k not 0 ,then no solution.

OpenStudy (anonymous):

wait so does it have infinite or no solutions

OpenStudy (anonymous):

2. y=-4x+a \[y=\frac{ 1 }{ 4 }x-a\] \[-4x+a=\frac{ 1 }{ 4 }x-a\] multiply by 4 -16x+4a=x-4a x=? then y=? only one solution in a giving different values to a we can get different solutions

OpenStudy (anonymous):

if k is 0 then infinite solutions if k is non zero then no solution.

OpenStudy (anonymous):

for A right

OpenStudy (anonymous):

i think your problem want solutions with slope, correct or not

OpenStudy (anonymous):

he just wants us to state how many solutions each system of equations have

OpenStudy (anonymous):

and why

OpenStudy (anonymous):

i put A into the calculatior and it gave the ordered pair (8a/17,-15a/17)

OpenStudy (anonymous):

2. the lines are perpendicular so cut in only one point so only one solution correct

OpenStudy (anonymous):

wait what about A? it has one solution?

OpenStudy (anonymous):

3.divide 1st eq. by 2 and 2nd by 3

OpenStudy (anonymous):

how can i divide one equation by the other?

OpenStudy (anonymous):

i am talking about 3rd question.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

nevermind this question has got me so confused

OpenStudy (anonymous):

thanks for your help tho

OpenStudy (anonymous):

x+3y=2c x+3y=2c so practically both are same and 3y=-x+c giving different values to x we get different solutions so it has infinite solutions.

OpenStudy (anonymous):

so a and b have only one solution and c has infinite?

OpenStudy (anonymous):

a. if k is non zero then no solution

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

only b has one solution.

OpenStudy (anonymous):

c has infinite solutions.

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

if k is non zero a has no solution.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

i am leaving now..Good night.

OpenStudy (anonymous):

take care@surjithayer